A 100mL solution 0.5 mol/L Sodium Chloride(NaCl) solution is combined with a 0.35 mol/L solution of Silver Nitrate(AgNO3). Write out the balanced equation and determine the volume of silver nitrate used.


1
Expert's answer
2021-04-12T00:04:34-0400

AgNO3(aq)+ NaCl(aq)→AgCl(s)+NaNO3(aq) 

Molar Mass of NaCl = 58.44g/mol

Molar Mass of AgNO3 = 169.87g/mol

= 58.44×0.5 = 29.22g

= 169.87 × 0.35 = 59.4545g

= 29.22+59.4545 = 88.6745

= (59.4545/88.6745)×100

= 67.048ml